# Oligonucleotide selective detection by levitated optomechanics

### v1

@@ Abstract
Figure 1: Functionaliaation: The process of functionalizing SiNPs with 25-mer deoxyadenosine monophosphate (25A) or 25-mer deoxythymidine monophosphate (25T) oligonucleotides, water and ZnCl2 solution.
@@
A physics explanation for the observed trap frequency shift for the oligonucleotide base-coated SiNP in the levitated optomechanical trap is described in Supplementary Section 1. This model, which ignores the ZnCl2 salt layer surrounding the surface of the SiNPs, finds that the polarizability-to-mass ratio $(\alpha/m)$ changes depending on the oligonucleotide. The estimates show that there is about a 1 kHz frequency shift per monolayer of DNA base molecule. There is a difference between adenine and thymine, and it is estimated that there is a frequency shift of 30 mHz corresponding to a single adenine molecule when compared to a single thymine base.
@@ Methods
Figure 6: Supplementary Fig. 1. The Omron MicroAIR U100 nebulizer. There are three main parts: a, is the interchangeable mesh piece where the aerosol droplets are released, b, is the reservoir where nanoparticle solutions are placed, and c, is the battery power source compartment.
@@
Previous optical trap experiments [15] used a setup similar to the one in this work. The iris was included such that when the aperture is reduced, the high-power Poisson spot is removed (Supplementary Fig. 2), decreasing the signal-to-noise ratio which increases the position resolution of the trapped particle that can be detected by the photodiode. This is necessary as the bright spot in the center of the beam overlaps the scattered field (Escat) and diverging field (Ediv). The bright Poisson spot forms from the flat mirror edges around the parabolic mirror and was used to confirm the laser alignment. The images in Supplementary Fig. 2 were captured by a Point Gray Research CMLN-13S2M-CS infrared camera.
Figure 7: Optical setup: A diagrammatic representation of the optical setup. with the laser system and optical beam path, focusing parabola in vacuum chamber and data collection by oscilloscope.
@@
Following the collection of raw data from the oscilloscope, the frequency plot of the time domain signal was obtained using a Fourier transform. This was carried out in Python code using the Optoanalysis package [22]. A Graphical User Interface (GUI) was developed on top of this code to enable seamless data entry as seen in Supplementary Fig. 4a. Following the display of the PSD, the recorded pressure was entered, as was the estimated peak frequency of the z, x and y degrees of motion, as well as a Yes/No value if the PSD Lorentzian curve was a good fit (Supplementary Fig. 4b-d). The data for each trapped particle was recorded as a row in a comma-separated value (CSV) table.
@@ Acknowledgements
Dielectric properties of oligonucleotides Typicaly, the static (frequency-independent) polarizabilities can be computed as response properties or finite field calculations. There have been extensive ab-initio density functional theory (DFT) calculations using various sets of density functionals [44, 45, 46] and complemented by Hartee-Fock methods [47], and classical electrodynamical calculations by Kramers-Kronig relations [48, 49].
Theoretical values for static polarizabilities can then be compared to experiments on dielectric properties of oligos for instance by Rayleigh scattering [50], by electrostatic force microscopy [51], or by optical extinction coefficient measurements [52]. Table 1 shows the values for $\alpha$ and m for ademine adn thymine used in this work.

### v2

@@ Abstract
Figure 1: Functionalization: The process of functionalizing SiNPs with 25-mer deoxyadenosine monophosphate (25A) or 25-mer deoxythymidine monophosphate (25T) oligonucleotides, water and ZnCl2 solution.
@@
A physics explanation for the observed trap frequency shift for the oligonucleotide base-coated SiNP in the levitated optomechanical trap is described in the Supplementary Information. This model, which ignores the ZnCl2 salt layer surrounding the surface of the SiNPs, finds that the polarizability-to-mass ratio $(\alpha/m)$ changes depending on the oligonucleotide. The estimates show that there is about a 1 kHz frequency shift per monolayer of DNA base molecule. There is a difference between adenine and thymine, and it is estimated that there is a frequency shift of 30 mHz corresponding to a single adenine molecule when compared to a single thymine base.
@@ Methods
Figure 6: The Omron MicroAIR U100 nebulizer. There are three main parts: a, is the interchangeable mesh piece where the aerosol droplets are released, b, is the reservoir where nanoparticle solutions are placed, and c, is the battery power source compartment.
@@
Previous optical trap experiments [15] used a setup similar to the one in this work. The iris was included such that when the aperture is reduced, the high-power Poisson spot is removed (Fig. 7), decreasing the signal-to-noise ratio which increases the position resolution of the trapped particle that can be detected by the photodiode. This is necessary as the bright spot in the center of the beam overlaps the scattered field (Escat) and diverging field (Ediv). The bright Poisson spot forms from the flat mirror edges around the parabolic mirror and was used to confirm the laser alignment. The images in Fig. 7 were captured by a Point Gray Research CMLN-13S2M-CS infrared camera.
Figure 7: IR images of decreasing iris apertures: A series of images demonstrating the removal of the Poisson spot with decreasing iris aperture. Closing the iris aperture decreases the size of the Poisson spot until the aperture size is equal to the diameter of the parabolic mirror.
@@
Figure 8: Optical setup: A diagrammatic representation of the optical setup with the laser system and optical beam path, focusing parabola in vacuum chamber and data collection by an oscilloscope.
@@
Following the collection of raw data from the oscilloscope, the frequency plot of the time domain signal was obtained using a Fourier transform. This was carried out in Python code using the Optoanalysis package [22]. A Graphical User Interface (GUI) was developed on top of this code to enable seamless data entry as seen in Fig. 9a. Following the display of the PSD, the recorded pressure was entered, as was the estimated peak frequency of the z, x and y degrees of motion, as well as a Yes/No value if the PSD Lorentzian curve was a good fit (Fig. 9b-d). The data for each trapped particle was recorded as a row in a comma-separated value (CSV) table.
@@ Acknowledgements
Dielectric properties of oligonucleotides Typically, the static (frequency-independent) polarizabilities can be computed as response properties or finite field calculations. There have been extensive ab-initio density functional theory (DFT) calculations using various sets of density functionals [44, 45, 46] and complemented by Hartee-Fock methods [47], and classical electrodynamics calculations by Kramers-Kronig relations [48, 49].
Theoretical values for static polarizabilities can then be compared to experiments on dielectric properties of oligos for instance by Rayleigh scattering [50], by electrostatic force microscopy [51], or by optical extinction coefficient measurements [52]. Table 1 shows the values for $\alpha$ and m for adenine and thymine used in this work.